A mass m2 = 15.0 kg is connected by a light cord to a mass m1 = 15.0 kg, which slides on a smooth horizontal surface. The pulley, of mass M = 1.00 kg, rotates about a frictionless axle and has a radius R = 0.200 m and a moment of inertia I = 0.0900 kg-m2 . The cord does not slip on the pulley.

a) What is the magnitude of the acceleration of m1?

b) What is the tension in the horizontal cord, T1?

c) What is the tension in the vertical cord, T2?
Answer #1

Hii

m2 0 for mass me mag- Tz = m, a 15x9.8 - T2 = 15xa for mass me Ti= mia=15a - for pulley T-T) XR = I a (T2-T ) XR = 0.ogx a @

Any query in solution then leave a comment

Please give me your valuable feedback

Thanks